# A transversality theorem for semi-algebraic sets with application to signal recovery from the second moment and cryo-EM

### v2

@@ 5 Permutation invariant separators for machine learning on sets
In this section, we consider the setting where $V=\mathbb{R}^{d\times n}$ is a representation of the permutation group $S_{n}$, with respect to the action of applying the same permutation to each of the $d$ rows of a matrix $X\in\mathbb{R}^{d\times n}$. This problem is motivated by machine learning applications for objects, like sets [51, 39] and graphs [50], which do not come with a natural ordering. In recent years, a line of work aims to design permutation invariant “separators” for $\mathbb{R}^{d\times n}$, or a subset $\mathcal{M}\subset\mathbb{R}^{d\times n}$ [19, 45, 3]. Namely, a permutation invariant mapping $F:\mathbb{R}^{d\times n}\to\mathbb{R}^{c}$ is separating on $\mathcal{M}$, if for all $X,Y\in\mathcal{M}$ we have that if $F(X)=F(Y)$ if and only if $X$ and $Y$ are related by an $S_{n}$ transformation. For efficiency purposes, it is desirable that the number $c$, the dimension of $F$ domain, is as small as possible. Moreover, to be useful for gradient descent type learning, the function $F$ is typically required to be differentiable, or at least continuous everywhere and piecewise differentiable.
When $d=1$, one possible solution is choosing $F_{1}(X)=\mathbf{sort}(X)$, which sorts the elements of the vector $X$ from small to large. This function is a permutation invariant separator and is continuous and piecewise linear.
When $d>1$, there are two common generalizations for the $d=1$ sorting operation. One is lexicographical sorting, where the ordering of $X\in\mathbb{R}^{d\times n}$ is determined by sorting the first row from small to large, in the case of ties sorting according to the second row, etc. Lexicographical sorting is a permutation invariant separator but is not continuous. The second generalization is row-wise sorting $F_{d}(X)=\mathbf{rowsort}(X)$, which is permutation invariant and continuous but is not separating. This can be easily seen by noting that $F_{d}(X)=F_{d}(Y)$ if and only if $Y$ is obtained from $X$ by applying $d$ different permutations independently to the $d$ different rows of $X\in\mathbb{R}^{d\times n}$. Thus, the “ambiguity group” of $F_{d}$ is not $S_{n}$ but the larger group $(S_{n})^{d}$. While $F_{d}$ is not separating on all of $\mathbb{R}^{d\times n}$, the situation may be different when considering a semialgebraic subset $\mathcal{M}\subseteq\mathbb{R}^{d\times n}$, since in this case the orbit of a given $X\in\mathcal{M}$ under the $(S_{n})^{d}$ symmetry may not have any intersections with $\mathcal{M}$ other than $X$ itself. Indeed, we can apply the second part of Theorem 2.2, using the fact that the orbit of the ambiguity group is finite and hence zero-dimensional, to obtain the following result.
@@ 6.1 The fiber lemma
We prove a general lemma, which is a generalization of [12, Lemma 2.1] and also resembles [23, Theorem 1.7 and Theorem 3.3].
Let $V$ be a representation of a compact group $H$ and let $\Theta$ be a semi-algebraic group of automorphisms of $V$. For each pair $x,y\in V$, set
@@
    If $K>M$, then for a generic vector $x\in\theta(\mathcal{M})$ if $y=h\cdot x\in\theta(\mathcal{M})$ for some $h\in H$, we must have $x\sim y$.
@@
    If $K>2M$ then for any vector $x\in\Theta(\mathcal{M})$ if $y=h\cdot x\in\Theta(\mathcal{M})$ for some $h\in H$, we must have $x\sim y$.
@@
$\mathcal{B}=\mathcal{B}(\mathcal{M},\Theta)=\{(x,y,\theta)\in\mathcal{M}\times\mathcal{M}\times\Theta\,|\,x\not\sim y\text{ but }\theta\cdot x\in H(\theta\cdot y)\}.$ (6.1)
@@
According to  [23, Lemma 1.8], we can bound the dimension of the incidence $\mathcal{B}$ by
@@
Now, if $K>2M$ then we obtain that $\dim(\mathcal{B})<\dim(\Theta)$. If $\phi$ is the projection $\phi(x,y,\theta)=\theta$, then $\dim(\phi(\mathcal{B}))<\dim(\Theta)$. Thus, a generic $\theta\in\Theta$ is not in $\phi(\mathcal{B})$. For all such $\theta$, if $x,y\in\theta\cdot\mathcal{M}$, then there exist $x^{\prime},y^{\prime}\in\mathcal{M}$ such that
$x=\theta\cdot x^{\prime},\text{ and }y=\theta\cdot y^{\prime}.$
Since $\theta\not\in\phi(B)$, we see that if $x\in Hy$ then $x^{\prime}\sim y^{\prime}$, which implies that $x\sim y$. This completes the proof of the second part of the lemma.
@@
$\mathcal{N}_{\theta}$ =$=\{x\in\mathcal{M}\,|\,\exists y\in\mathcal{M}\text{ such that }(x,y,\theta)\in\mathcal{B}\}$ (6.3)
=$=\{x\in\mathcal{M}\,|\,\exists y\in\mathcal{M}\text{ such that }x\not\sim y\text{ but }(\theta\cdot x)\in H(\theta\cdot y)\},$
@@
We now prove that the generic fiber of $\phi\colon\mathcal{B}\to\Theta$ has dimension less than $M$. By [18, Proposition 2.9.10], the semi-algebraic set $\Theta$ is a disjoint union of Nash submanifolds $\Theta_{1},\ldots,\Theta_{s}$. Clearly, it suffices to prove that $\dim(\phi^{-1}(\theta)\cap\mathcal{B})<M$ for a generic $\theta$ in any Nash submanifold $\Theta_{\ell}$, where $\dim\Theta_{\ell}=\dim\Theta$. With this observation, we can reduce to the case that $\Theta$ is a manifold.
Now we decompose $\mathcal{B}$ as a disjoint union of Nash submanifolds $\bigcup_{\ell=1}^{L}U_{\ell}$ and let $\mathcal{L}\subseteq\{1,\ldots,L\}$ denote the subset of indices $\ell$ for which $\phi(U_{\ell})$ attains its maximal value $\dim(\Theta)$. For every fixed $\ell\in\mathcal{L}$, the semi-algebraic version of Sard’s Theorem [18, Theorem 9.6.2.] implies that the generic $\theta\in\phi(U_{\ell})$ is a regular value (i.e., not the image of a critical point) of the restriction of $\phi$ to $U_{\ell}$. By the pre-image theorem [47, Theorem 9.9], every regular value $\theta$ is either not in the image of $\phi|_{U_{\ell}}$, or we have the equality
@@
The semi-algebraic set which is the union of the images of the critical points of $\{\phi_{\ell}\}_{\ell\in\mathcal{L}}$ and $\cup_{\ell\not\in\mathcal{L}}\phi(U_{\ell})$ has dimension strictly less than $\dim\Theta$. If $\theta\in\Theta$ lies in the complement of this set, then we have that the fiber
@@
is at most $\dim\mathrm{GL}(V)-K$. If $x$ and $y$ are parallel, then $Ax$ and $Ay$ are parallel for all $A\in\mathrm{GL}(V)$. Since $H$ acts by unitary transformations, $Ax$ is in the $H$-orbit of $Ay$ if and only if $|Ax|=|Ay|$ or equivalently if and only if $|x|=|y|$. Since $x$ and $y$ are parallel, this means $x=\pm y$, so $\operatorname{GL}(x,y)$ is empty unless the vectors $x$ and $y$ are $\mathrm{GL}(V)$-equivalent.
Now, suppose that $x$ and $y$ are linearly independent. Then we can choose an ordered basis $b_{1},b_{2},\ldots,b_{N}$ with $b_{1}=x$ and $b_{2}=y$. If we express an element $A\in\mathrm{GL}(V)$ as a matrix with respect to this ordered basis, then $Ax$ is the first row of $A$ and $Ay$ is the second row of $A$. The condition that $Ax$ lies in the $H$-orbit of $Ay$ implies that the first row of the matrix $A$ lies in a real algebraic subset of $\mathbb{R}^{N}$ of dimension at most $k(H)$. Hence,
@@
$S_{1}$ =$=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T=-(y+T)\}$
$S_{2}$ =$=\{(L,T)\in\mathrm{Aff}(x,y)\quad x+T\neq-(y+T)\text{ but they are linearly dependent}\}$
$S_{3}$ =$=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T\text{ and }y+T\text{ are linearly independent}\}.$
@@
We start with $S_{1}$. If $(L,T)$ is in $S_{1}$, then there is a unique $T\in V$ such that $x+T=-(y+T)$, namely $T=-\frac{1}{2}(x+y)$. Therefore, the dimension of $S_{1}$ is at most the dimension of $\mathrm{GL}(V)$, which gives us a bound on the dimension,
$\dim(S_{1})=\dim(\mathrm{GL}(V))=\dim(\mathrm{Aff}(V))-\dim(V)\leq\dim(\mathrm{Aff}V)-K,$
@@
Next, we prove that $S_{2}$ is empty. If $(L,T)$ is in $S_{2}$, then $x+T$ and $y+T$ are linearly dependent, but $x+T\neq-(y+T)$ by the definition of $S_{2}$. Also $x+T\neq y+T$ since $x\neq y$. In particular, $x+T$ and $-(y+T)$ cannot both be zero. Assume without loss of generality that $x+T\neq 0$, then there exists some $\lambda\in\mathbb{R}$ such that $|\lambda|\neq 1$ and $y+T=\lambda(x+T)$. If $L(x+T))=h(L(y+T))$ for some $h\in H$, then in particular $L(x+T)$ and $L(y+T)=\lambda L(x+T)$ have the same norm because we assume that $H$ acts by orthogonal transformations. But this is impossible since we assume $|\lambda|\neq 1$.
@@
Once again, it is sufficient to bound the dimension of $\mathrm{O}(x,y)=\{A\in\mathrm{O}(V)\,|\,Ax\in H(Ay)\}$. If $x$ and $y$ are linearly dependent, then $\mathrm{O}(x,y)$ is empty unless $x=\pm y$ and then $x\sim y$. On the other hand, if $x,y$ are linearly independent, then after normalizing we can assume that $x$ is a unit vector. Choose an orthonormal basis $e_{1},\ldots,e_{N}$ for $V$ such that $e_{1}=x$ and $y=ae_{1}+be_{2}$ for some constants $a,b$ with $b\neq 0$ (because $y$ is not parallel to $x$). Now suppose that $A\in\mathrm{O}(V)$ is an orthogonal matrix. The condition that $Ax=hAy$ for some $h\in H$ is equivalent to the condition that $Ay=aAe_{1}+bAe_{2}$ is in the $H$-orbit of $Ae_{1}$. The vectors $Ae_{1}$ and $Ae_{2}$ are just the first two rows of the matrix $A$. Call these vectors $w_{1}$ and $w_{2}$ respectively.
For a given choice of first row $w_{1}$, there is an $(n-2)$-dimensional set of possible vectors $w_{2}$, which can be the second row of the orthogonal matrix $A$. The condition that $Ay\in HAx$ means that we must choose $w_{2}$ to be in the real algebraic subset $-ab^{-1}w_{1}+b^{-1}Hw_{1}\subset\mathbb{R}^{N}$, which has dimension at most $k(H)$. Thus, $w_{2}$ is taken from a subset of $S^{n-2}$ of codimension at least $n-2-k(H)$, which is equal to $K-2$. On the other hand, once we have selected the rows $w_{1},w_{2}$, we impose no additional conditions on the subsequent rows of $A$. Hence, $\dim A(x,y)\leq\dim\mathrm{O}(V)-K+2$.
Now, suppose that the orbits of $H$ are connected. First, note that $\dim Hw_{1}\neq 0$. If this were the case, then $Hw_{1}=w_{1}$, so $w_{2}=-ab^{-1}w_{1}+b^{-1}w_{1}$. However, this is impossible because $w_{2}$ and $w_{1}$ are linearly independent. Hence, we may suppose that $Hw_{1}$ is a connected real algebraic subset of $\mathbb{R}^{N}$ of positive dimension. Because orbits are also submanifolds, the connectedness of $Hw_{1}$ implies that it is also irreducible as an algebraic set. Since $w_{2}=-ab^{-1}w_{1}+b^{-1}hw_{1}$ is a unit vector, we impose the additional condition that
@@
Since $Hw_{1}$ is positive dimensional, it contains at least one vector $hw_{1}\neq\pm w_{1}$, so the algebraic function
@@
is not constant on the irreducible algebraic set $Hw_{1}$ unless $a=0$ and $b=\pm 1$. Thus, if $x,y$ are not orthogonal vectors with the same magnitude, then the condition that $w_{2}$ is a unit vector shows that $w_{2}$ is taken from an algebraic subset of $S^{n-2}$ of dimension at most $k(H)-1$, so in this case $\dim\mathrm{O}(x,y)\leq\dim\mathrm{O}(V)-K+1$.
On the other hand, if $x,y$ are orthogonal vectors of the same magnitude, then after normalizing we must have that $w_{1}=x$ and $w_{2}=y$. The requirement that $\langle w_{1},w_{2}\rangle=0$ imposes the condition that $\langle w_{1},hw_{1}\rangle=0$. This condition cannot be identically satisfied on the orbit $Hw_{1}$, since it is not satisfied at $w_{1}\in Hw_{1}$. Thus, once again we see that $\dim\mathrm{O}(x,y)\leq\dim\mathrm{O}(V)-K+1$.

### v3

@@ 5 Permutation invariant separators for machine learning on sets
In this section, we consider the setting where $V=\mathbb{R}^{d\times n}$ is a representation of the permutation group $S_{n}$, with respect to the action of applying the same permutation to each of the $d$ rows of a matrix $X\in\mathbb{R}^{d\times n}$. This problem is motivated by machine learning applications for objects, like sets [56, 41] and graphs [55], which do not come with a natural ordering. In recent years, a line of work aims to design permutation invariant “separators” for $\mathbb{R}^{d\times n}$, or a subset $\mathcal{M}\subset\mathbb{R}^{d\times n}$ [20, 50, 3]. Namely, a permutation invariant mapping $F:\mathbb{R}^{d\times n}\to\mathbb{R}^{c}$ is separating on $\mathcal{M}$, if for all $X,Y\in\mathcal{M}$ we have that $F(X)=F(Y)$ if and only if $X$ and $Y$ are related by an $S_{n}$ transformation. For efficiency purposes, it is desirable that the number $c$, the dimension of the codomain of $F$, is as small as possible. Moreover, to be useful for gradient descent-type learning, the function $F$ is typically required to be differentiable, or at least continuous everywhere and piecewise differentiable.
When $d=1$, one possible solution is choosing $F_{1}(X)=\mathbf{sort}(X)$, which sorts the elements of the vector $X$ from smallest to largest. This function is a permutation invariant separator and is continuous and piecewise linear.
When $d>1$, there are two common generalizations for the $d=1$ sorting operation. One is lexicographical sorting, where the ordering of $X\in\mathbb{R}^{d\times n}$ is determined by sorting the first row from small to large, in the case of ties, sorting according to the second row, etc. Lexicographical sorting is a permutation invariant separator, but is not continuous. The second generalization is sorting each row independently, which we denote by $F_{d}(X)=\mathbf{rowsort}(X)$. This is permutation invariant and continuous, but is not separating. This can be easily seen by noting that $F_{d}(X)=F_{d}(Y)$ if and only if $Y$ is obtained from $X$ by applying $d$ different permutations independently to the $d$ different rows of $X\in\mathbb{R}^{d\times n}$. Thus, the “ambiguity group” of $F_{d}$ is not $S_{n}$ but the larger group $(S_{n})^{d}$. While $F_{d}$ is not separating on all of $\mathbb{R}^{d\times n}$, the situation may be different when considering a semialgebraic subset $\mathcal{M}\subseteq\mathbb{R}^{d\times n}$, since in this case the orbit of a given $X\in\mathcal{M}$ under the $(S_{n})^{d}$ symmetry may not have any intersections with $\mathcal{M}$ other than $X$ itself. Indeed, we can apply the second part of Theorem 2.2, using the fact that the orbit of the ambiguity group is finite and hence zero-dimensional, to obtain the following result.
@@ 6.1 The fiber lemma
We prove a general lemma, which is a generalization of [13, Lemma 2.1] and also resembles [24, Theorem 1.7 and Theorem 3.3].
Let $V$ be an orthogonal representation of a compact Lie group $H$ and let $\Theta$ be a semi-algebraic group of automorphisms of $V$. For each pair $x,y\in V$, set
@@
    If $K>M$, then for a generic vector $z\in\theta\cdot\mathcal{M}$ if there is a vector $w\in\theta\cdot\mathcal{M}$ such that $z=h\cdot w$ for some $h\in H$, we must have $z\sim w$.
@@
    If $K>2M$ then for any vector $z\in\theta\cdot\mathcal{M}$ if there is a vector $w\in\theta\cdot\mathcal{M}$ such that $z=h\cdot w$ for some $h\in H$, we must have $z\sim w$.
Remark 6.2.
We note that Lemma 6.1 holds for any integer $K$ that satisfies the inequalities
$\dim(\Theta(x,y))\leq\dim(\Theta)-K,\quad\forall x,y\in\mathcal{M},\text{ with }x\not\sim y$
and $K>M$ (resp. $K>2M$). However, when we apply the lemma, we always take $K=\dim V-k(H)$, where $k(H)$ is the maximal dimension of an $H$-orbit, as defined in (2.1).
@@
$\mathcal{B}=\mathcal{B}(\mathcal{M},\Theta)=\{(x,y,\theta)\in\mathcal{M}\times\mathcal{M}\times\Theta\,|\,x\not\sim y\text{ {\color[rgb]{0,0,0} and} }\theta\cdot x\in H(\theta\cdot y)\}.$ (6.1)
@@
According to  [24, Lemma 1.8], we can bound the dimension of the incidence $\mathcal{B}$ by
@@
If $K>2M$, then (6.2) implies that $\dim(\mathcal{B})<\dim(\Theta)$. In particular, if $\phi$ is the projection $\phi(x,y,\theta)=\theta$, then $\dim(\phi(\mathcal{B}))<\dim(\Theta)$. It follows that a generic $\theta\in\Theta$ is not in $\phi(\mathcal{B})$. Suppose that $z,w\in\theta\cdot\mathcal{M}$ with $z=\theta\cdot x$ and $w=\theta\cdot y$ for some $x,y\in\mathcal{M}$. If $\theta\not\in\phi(B)$, then by definition of $\mathcal{B}$ we know that if $z=h\cdot w$ for some $h\in H$ then $x\sim y$. Thus, $z=\theta\cdot x$ and $w=\theta\cdot y$ are also equivalent, which completes the proof of the second part of the lemma.
@@
$\mathcal{N}_{\theta}$ $=\{x\in\mathcal{M}\,|\,\exists y\in\mathcal{M}\text{ such that }(x,y,\theta)\in\mathcal{B}\}$ (6.3)
$=\{x\in\mathcal{M}\,|\,\exists y\in\mathcal{M}\text{ such that }x\not\sim y\text{ but }(\theta\cdot x)\in H(\theta\cdot y)\},$
@@
We now prove that the generic fiber of $\phi\colon\mathcal{B}\to\Theta$ has dimension less than $M$. By [19, Proposition 2.9.10], the semi-algebraic set $\Theta$ is a disjoint union of Nash submanifolds $\Theta_{1},\ldots,\Theta_{s}$. Clearly, it suffices to prove that $\dim(\phi^{-1}(\theta)\cap\mathcal{B})<M$ for a generic $\theta$ in any Nash submanifold $\Theta_{\ell}$, where $\dim\Theta_{\ell}=\dim\Theta$. With this observation, we can reduce to the case that $\Theta$ is a manifold.
Now we decompose $\mathcal{B}$ as a disjoint union of Nash submanifolds777 A semi-algebraic subset of $A\subset\mathbb{R}^{N}$ is called a Nash submanifold of dimension $d$ if, for every point $x\in A$, there is an open neighborhood $x\in U\subset\mathbb{R}^{N}$ and a $C^{\infty}$ differentiable function $\varphi\colon U\to\mathbb{R}^{d}$ such that the restriction of $\varphi$ to $U\cap A$ is a diffeomorphism [19, Definition 2.9.9]. $\bigcup_{\ell=1}^{L}U_{\ell}$ and let $\mathcal{L}\subseteq\{1,\ldots,L\}$ denote the subset of indices $\ell$ for which $\phi(U_{\ell})$ attains its maximal value $\dim(\Theta)$. For every fixed $\ell\in\mathcal{L}$, the semi-algebraic version of Sard’s Theorem [19, Theorem 9.6.2.] implies that the generic $\theta\in\phi(U_{\ell})$ is a regular value (i.e., not the image of a critical point) of the restriction of $\phi$ to $U_{\ell}$. By the pre-image theorem [52, Theorem 9.9], every regular value $\theta$ is either not in the image of $\phi|_{U_{\ell}}$, or we have the equality
@@
The semi-algebraic set, which is the union of the images of the critical points of $\{\phi_{\ell}\}_{\ell\in\mathcal{L}}$ and $\cup_{\ell\not\in\mathcal{L}}\phi(U_{\ell})$ has dimension strictly less than $\dim\Theta$. If $\theta\in\Theta$ lies in the complement of this set, then we have that the fiber
@@
is at most $\dim\mathrm{GL}(V)-K$. If $x$ and $y$ are parallel, then $Ax$ and $Ay$ are parallel for all $A\in\mathrm{GL}(V)$. Since $H$ acts by unitary transformations, $Ax$ is in the $H$-orbit of $Ay$ only if $|Ax|=|Ay|$ or equivalently only if $|x|=|y|$. (Here $|\cdot|$ denotes the usual Euclidean norm.) Since $x$ and $y$ are parallel, this means $x=\pm y$, so $\operatorname{GL}(x,y)$ is empty unless the vectors $x$ and $y$ are $\mathrm{GL}(V)$-equivalent.
Now, suppose that $x$ and $y$ are linearly independent. Then we can choose an ordered basis $b_{1},b_{2},\ldots,b_{N}$ with $b_{1}=x$ and $b_{2}=y$. If we express an element $A\in\mathrm{GL}(V)$ as a matrix with respect to this ordered basis, then $Ax$ is the first column of $A$ and $Ay$ is the second column of $A$. The condition that $Ax$ lies in the $H$-orbit of $Ay$ implies that the first column of the matrix $A$ lies in a real algebraic subset of $\mathbb{R}^{N}$ of dimension at most $k(H)$. Hence,
@@
$S_{1}$ $=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T=-(y+T)\}$
$S_{2}$ $=\{(L,T)\in\mathrm{Aff}(x,y)\quad x+T\neq-(y+T)\text{ but they are linearly dependent}\}$
$S_{3}$ $=\{(L,T)\in\mathrm{Aff}(x,y)|\quad x+T\text{ and }y+T\text{ are linearly independent}\}.$
@@
We start with $S_{1}$. If $(L,T)$ is in $S_{1}$, then there is a unique $T\in V$ such that x+T=(y+T)x+T=-(y+T), namely $T=-\frac{1}{2}(x+y)$. Therefore, the dimension of $S_{1}$ is at most the dimension of $\mathrm{GL}(V)$, which gives us a bound on the dimension,
$\dim(S_{1})=\dim(\mathrm{GL}(V))=\dim(\mathrm{Aff}(V))-\dim(V)\leq\dim(\mathrm{Aff}(V))-K,$
@@
Next, we prove that $S_{2}$ is empty. If $(L,T)$ is in $S_{2}$, then $x+T$ and $y+T$ are linearly dependent, but $x+T\neq-(y+T)$ by the definition of $S_{2}$. Also $x+T\neq y+T$ since $x\neq y$. In particular, $x+T$ and (y+T)-(y+T) cannot both be zero. Assume without loss of generality that $x+T\neq 0$, then there exists some $\lambda\in\mathbb{R}$ such that $|\lambda|\neq 1$ and $y+T=\lambda(x+T)$. If $L(x+T)=h(L(y+T))$ for some $h\in H$, then in particular $L(x+T)$ and $L(y+T)=\lambda L(x+T)$ have the same norm because we assume that $H$ acts by orthogonal transformations. But this is impossible since we assume $|\lambda|\neq 1$.
@@
Once again, it is sufficient to bound the dimension of $\mathrm{O}(x,y)=\{A\in\mathrm{O}(V)\,|\,Ax\in H(Ay)\}$. If $x$ and $y$ are linearly dependent, then $\mathrm{O}(x,y)$ is empty unless $x=\pm y$ and then $x\sim y$. On the other hand, if $x,y$ are linearly independent, then after normalizing, we can assume that $x$ is a unit vector. Choose an orthonormal basis $e_{1},\ldots,e_{N}$ for $V$ such that $e_{1}=x$ and $y=ae_{1}+be_{2}$ for some constants $a,b$ with $b\neq 0$ (because $y$ is not parallel to $x$). Now suppose that $A\in\mathrm{O}(V)$ is an orthogonal matrix. The condition that $Ax=hAy$ for some $h\in H$ is equivalent to the condition that $Ay=aAe_{1}+bAe_{2}$ is in the $H$-orbit of $Ae_{1}$. The vectors $Ae_{1}$ and $Ae_{2}$ are just the first two columns of the matrix $A$. Call these vectors $w_{1}$ and $w_{2}$ respectively.
For a given choice of first column $w_{1}$, there is an (n2)(n-2)-dimensional set of possible vectors $w_{2}$, which can be the second column of the orthogonal matrix $A$. The condition that $Ay\in HAx$ means that we must choose $w_{2}$ to be in the real algebraic subset $ab^{-1}w_{1}+b^{-1}Hw_{1}\subset\mathbb{R}^{N}$, which has dimension at most $k(H)$. Thus, $w_{2}$ is taken from a subset of $S^{n-2}$ of codimension at least n2k(H)n-2-k(H), which is equal to K2K-2. On the other hand, once we have selected the columns $w_{1},w_{2}$, we impose no additional conditions on the subsequent columns of $A$. Hence, $\dim A(x,y)\leq\dim\mathrm{O}(V)-K+2$.
Now, suppose that the orbits of $H$ are connected. First, note that if $\dim Hw_{1}=0$ then $Ay=Ax$ so $x$ and $y$ are linear dependent and $O(x,y)$ is empty. Hence, we may suppose that $Hw_{1}$ is a connected real algebraic subset of $\mathbb{R}^{N}$ of positive dimension. Because orbits are also smooth submanifolds, the connectedness of $Hw_{1}$ implies that it is also irreducible as an algebraic set. Since $w_{2}=-ab^{-1}w_{1}+b^{-1}hw_{1}$ is a unit vector, we impose the additional condition that
@@
Since $Hw_{1}$ is positive-dimensional, it contains at least one vector $hw_{1}\neq\pm w_{1}$, so the algebraic function
@@
is not constant on the irreducible algebraic set $Hw_{1}$ unless $a=0$ and $b=\pm 1$. Thus, if $x,y$ are not orthogonal vectors with the same magnitude, then the condition that $w_{2}$ is a unit vector shows that $w_{2}$ is taken from an algebraic subset of $S^{n-2}$ of dimension at most k(H)1k(H)-1, so in this case $\dim\mathrm{O}(x,y)\leq\dim\mathrm{O}(V)-K+1$.
On the other hand, if $x,y$ are orthogonal vectors of the same magnitude, then after normalizing we must have that $w_{1}=x$ and $w_{2}=\pm y$. The requirement that $\langle w_{1},w_{2}\rangle=0$ imposes the condition that $\langle w_{1},hw_{1}\rangle=0$. This condition cannot be identically satisfied on the orbit $Hw_{1}$, since it is not satisfied at $w_{1}\in Hw_{1}$. Thus, once again we see that $\dim\mathrm{O}(x,y)\leq\dim\mathrm{O}(V)-K+1$.